# Preference-based Reinforcement Learning beyond Pairwise Comparisons: Benefits of Multiple Options

### v1

@@ Abstract
We study online preference-based reinforcement learning (PbRL) with the goal of improving sample efficiency. While a growing body of theoretical work has emerged—motivated by PbRL’s recent empirical success, particularly in aligning large language models (LLMs)—most existing studies focus only on pairwise comparisons. A few recent works [96, 51, 79] have explored using multiple comparisons and ranking feedback, but their performance guarantees fail to improve—and can even deteriorate—as the feedback length increases, despite the richer information available. To address this gap, we adopt the Plackett–Luce (PL) model for ranking feedback over action subsets and propose M-AUPO, an algorithm that selects multiple actions by maximizing the average uncertainty within the offered subset. We prove that M-AUPO achieves a suboptimality gap of $\tilde{\mathcal{O}}\!\left(\frac{d}{T}\!\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\right)$, where $T$ is the total number of rounds, $d$ is the feature dimension, and $|S_{t}|$ is the size of the subset at round $t$. This result shows that larger subsets directly lead to improved performance and, notably, the bound avoids the exponential dependence on the unknown parameter’s norm, which was a fundamental limitation in most previous works. Moreover, we establish a near-matching lower bound of $\Omega\Big(\frac{d}{K\sqrt{T}}\Big)$, where $K$ is the maximum subset size. To the best of our knowledge, this is the first theoretical result in PbRL with ranking feedback that explicitly shows improved sample efficiency as a function of the subset size.
@@ 1 Introduction
The framework of Preference-based Reinforcement Learning (PbRL) [12, 83, 84, 72] was introduced to address the difficulty of designing effective reward functions, which often demands substantial and complex engineering effort [82, 84]. PbRL has been successfully applied in diverse domains, including robot training, stock prediction, recommender systems, and clinical trials [30, 67, 18, 38, 54]. Notably, PbRL also serves as a foundational framework for Reinforcement Learning from Human Feedback (RLHF) when feedback is provided in the form of preferences rather than explicit scalar rewards. This preference-based approach has proven highly effective in aligning Large Language Models (LLMs) with human values and preferences [18, 59, 64].
@@
Given its practical success, the field has also seen significant theoretical advances [16, 49, 72, 96, 89, 94, 93, 86, 74, 53, 13, 66, 22, 19, 51, 77, 73, 79, 88, 14, 39]. However, despite this progress, most existing models remain limited to handling only pairwise comparison feedback. A few works [96, 51, 79] explore the more general setting of multiple comparisons, offering a strict extension beyond the pairwise case. Zhu et al. [96] study the offline setting, where a dataset of questions (or contexts) along with corresponding ranking feedback over $K$ answers (or actions), labeled by human annotators, is available. Mukherjee et al. [51] investigate the online learning-to-rank problem [63], where a dataset of questions with $K$ candidate answers is provided, but no feedback is initially available. Thekumparampil et al. [79] consider a context-free setting (i.e., a singleton context), and the goal is to learn the ranking of $N\geq K$ answers based on ranking feedback obtained from subsets of size $K$. However, all of their theoretical performance guarantees fail to show that using multiple comparisons provides any advantage over the pairwise setting (see Table 1). This is counterintuitive, as ranking feedback is inherently more informative than pairwise feedback. Specifically, since a ranking over $K$ actions provides $\binom{K}{2}$ pairwise comparisons, it should, in principle, enable faster learning and lead to stronger performance guarantees.
@@
Table 1: Comparisons of settings and theoretical guarantees in related works on PbRL with ranking feedback. Here, $T$ denotes the number of rounds (or the number of data points in the offline setting), $K$ is the (maximum) size of the offered action set (i.e., assortment), and $d$ is the feature dimension, and $1/\kappa=\mathcal{O}(e^{B})$. $\rho$ represents the unknown context distribution. Here, $\tilde{\mathcal{O}}$ hides logarithmic factors and polynomial dependencies on $B$. “Sq. Pred. Error” refers to the squared prediction error.
@@
Zhu et al. [96] Offline Accessible $\mathcal{X}$ Given Suboptimality $\tilde{\mathcal{O}}\left(\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K^{2}}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\kappa}}\sqrt{\frac{d}{T}}\right)$
Mukherjee et al. [51] Online Accessible $\mathcal{X}$ Given Pred. Error $\tilde{\mathcal{O}}\left(\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K^{3}}d}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\kappa}\sqrt{T}}\right)$
Thekumparampil et al. [79] Online No context Select $K$ Pred. Error $\tilde{\mathcal{O}}\left(\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K^{3}}d}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\kappa}\sqrt{T}}\right)$
@@
In this paper, we assume that the ranking feedback follows the Plackett-Luce (PL) model [62, 47], where, in each round, the learner receives ranking feedback over a subset of up to $K$ actions (with $K\leq N$) selected from a universe of $N$ actions. This problem setup is closely related to that of Thekumparampil et al. [79]; however, unlike their work, which focuses solely on a context-free setting (or equivalently, a fixed singleton context), we study a more general setting where contexts are diverse and drawn from an unknown distribution.
@@
Furthermore, our suboptimality gap eliminates the exponential dependence on the parameter norm bound, $\mathcal{O}(e^{B})$, in the leading term, by employing novel matrix concentration inequalities for the Hessian matrix $H_{t}$ (see the proof sketch in Section 5.1 for details). This represents a significant improvement over most prior works, where performance guarantees typically depend on $\mathcal{O}(e^{B})$ [68, 72, 96, 89, 94, 19, 88, 79, 39]. Very recently, a few works [14, 20] have successfully avoided the $\mathcal{O}(e^{B})$ dependency by relying on auxiliary techniques or additional information—such as specialized sampling schemes [14] or prior knowledge of $\kappa$ [20]—which, however, are often impractical. Moreover, their methods are limited to pairwise comparison settings. In contrast, our approach eliminates the $\mathcal{O}(e^{B})$ dependency without using any auxiliary techniques and considers more general ranking feedback beyond pairwise comparisons. Our main contributions are summarized as follows:
@@
  • Free of $\mathcal{O}(e^{B})$ dependency: Our result eliminates the exponential dependence on the parameter norm bound, $\mathcal{O}(e^{B})$, in the leading term, without relying on any auxiliary techniques. This demonstrates that the $\mathcal{O}(e^{B})$ dependence commonly observed in PbRL (or RLHF) and dueling bandit analyses is not fundamentally necessary, but rather an artifact of loose analysis. We believe that our key lemmas (Lemmas 1 and E.1) can be directly applied to existing PbRL or dueling bandit analyses—including regret-minimization settings—whenever elliptical potential lemmas are used, without requiring any modification to the original algorithm. To the best of our knowledge, this is the first PbRL work with ranking feedback involving more than two options that avoids $\mathcal{O}(e^{B})$ dependence.
  • Lower bound: We establish a near-matching lower bound of $\Omega\!\left(\frac{d}{K\sqrt{T}}\right)$ under the Plackett–Luce (PL) model with ranking feedback, matching our upper bound up to a factor of $K$. This result demonstrates that incorporating richer ranking information (i.e., larger $K$) provably enhances sample efficiency.
@@ 2 Related Works
Fueled by the remarkable success of LLMs [18, 59, 64], the theoretical study of PbRL has rapidly emerged as a central focus within the research community. Early work in this area traces back to the dueling bandits literature [91, 98, 70, 8].
@@
Dueling bandits. The dueling bandit framework, introduced by Yue et al. [91], departs from the classical multi-armed bandit setting by requiring the learner to select two arms and observe only their pairwise preference. For general preferences, a single best arm that is globally dominant may not exist. To address this, various alternative winners have been proposed, including the Condorcet winner [97, 36], Copeland winner [98, 85, 37], Borda winner [31, 25, 28, 71, 87], and von Neumann winner [65, 24, 7], each with its own corresponding performance metric.
@@
To address scalability and contextual information, Saha [68] proposed a structured contextual dueling bandit setting in which preferences are modeled using a Bradley–Terry–Luce (BTL) model [11] based on the unknown intrinsic rewards of each arm. In a similar setting, Bengs et al. [9] studied a contextual linear stochastic transitivity model, and Di et al. [21] proposed a layered algorithm that achieves variance-aware regret bounds. However, most prior dueling bandit works suffer from an exponential dependency of $\mathcal{O}(e^{B})$. In recent work, only a few studies [20, 14] have succeeded in eliminating the $\mathcal{O}(e^{B})$ dependency by incorporating additional complex subroutines.
@@
Preference-based reinforcement learning (PbRL). Building upon this line of work, subsequent research has extended the dueling bandit framework to the RL framework, considering both online [90, 54, 16, 72, 86] and offline settings [96, 94, 46]. More recently, under the active learning framework—where the full set of contexts $\mathcal{X}$ is accessible—many studies aim to improve sample efficiency by selecting prompts either based on the differences in estimated rewards for their responses [52] or through D-optimal design methods [49, 73, 19, 51, 79, 39]. However, most of these works focus exclusively on pairwise preference feedback and cannot be extended to more general ranking feedback cases. Mukherjee et al. [51] study the online learning-to-rank problem when prompts are given along with $K$ candidate answers, while Thekumparampil et al. [79] investigate learning to rank $N\geq K$ answers from partial rankings over $K$ answers, but under a context-free setting. In this paper, we consider a stochastic contextual setting (more general than Thekumparampil et al. [79]), where contexts are sampled from an unknown but fixed distribution, and aim to minimize the suboptimality gap using ranking feedback of up to length $K$.
For further related work, see Appendix A.
@@ 3 Problem Setting and Preliminaries
Let ${\mathcal{S}}:=\{S\subseteq\mathcal{A}\mid 2\leq|S|\leq K\}$ be the collection of all action subsets whose sizes range from $2$ to $K$. For any $S\in{\mathcal{S}}$, let $\sigma$ denote the labeler’s ranking feedback—that is, a permutation of the elements in $S$. We write $\sigma_{j}$ for the $j$-th most preferred action under $\sigma$. We model the distribution of such rankings using the Plackett-Luce (PL) model [62, 47], defined as:
@@
Here, $r_{\bm{\theta}^{\star}}$ represents a reward model parameterized by the unknown parameter $\theta^{\star}$.
When K=2K\!=\!2, this reduces to the pairwise comparison framework considered in the Bradley-Terry-Luce (BTL) model [11]. The probability that $a$ is preferred to $a^{\prime}$ given $x$ can be expressed as:
@@
Let $\phi:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ be a known feature map satisfying $\max_{x,a}\|\phi(x,a)\|_{2}\leq 1$, and let $\bm{\theta}^{\star}\in\mathbb{R}^{d}$ denote the true but unknown parameter. The reward is assumed to follow a linear structure given by $r_{\bm{\theta}^{\star}}(x,a)=\phi(x,a)^{\top}\bm{\theta}^{\star}$. To ensure identifiability of $\bm{\theta}^{\star}$, we assume that $\bm{\theta}^{\star}\in\Theta$, where $\Theta:=\{\bm{\theta}\in\mathbb{R}^{d}\mid\|\bm{\theta}\|_{2}\leq B\}$. Without loss of generality, we assume $B\geq 1$.
At each round $t\in[T]$, a context $x_{t}\in\mathcal{X}$ is drawn from a fixed but unknown distribution $\rho$. Given the context $x_{t}$, the learning agent selects a subset of actions $S_{t}\in{\mathcal{S}}$—referred to as an assortment throughout the paper—and receives a ranking over $S_{t}$ as feedback, generated according to the PL model. After $T$ rounds of interaction with the labeler, the goal is to output a policy $\widehat{\pi}_{T}:\mathcal{X}\rightarrow\mathcal{A}$ that minimizes the suboptimality gap, defined as:
$\textup{{SubOpt}}(T):=\mathbb{E}_{x\sim\rho}\left[r_{\bm{\theta}^{\star}}\left(x,\pi^{\star}(x)\right)-r_{\bm{\theta}^{\star}}\left(x,\widehat{\pi}_{T}(x)\right)\right],$
where $\pi^{\star}(x)=\mathop{\mathrm{argmax}}_{a}r_{\bm{\theta}^{\star}}(x,a)$ is the optimal policy under the true reward $r_{\bm{\theta}^{\star}}$.
3.1 Loss Functions and Rank Breaking

### v2

@@ Abstract
We study online preference-based reinforcement learning (PbRL) with the goal of improving sample efficiency. While a growing body of theoretical work has emerged—motivated by PbRL’s recent empirical success, particularly in aligning large language models (LLMs)—most existing studies focus only on pairwise comparisons. A few recent works [zhu2023principled, mukherjee2024optimal, thekumparampil2024comparing] have explored using multiple comparisons and ranking feedback, but their performance guarantees fail to improve—and can even deteriorate—as the feedback length increases, despite the richer information available. To address this gap, we adopt the Plackett–Luce (PL) model for ranking feedback over action subsets and propose M-AUPO, an algorithm that selects multiple actions by maximizing the average uncertainty within the offered subset. We prove that M-AUPO achieves a suboptimality gap of $\tilde{\mathcal{O}}\!\left(\frac{d}{T}\!\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\right)$, where $T$ is the total number of rounds, $d$ is the feature dimension, and $|S_{t}|$ is the size of the subset at round $t$. This result shows that larger subsets directly lead to improved performance and, notably, the bound avoids the exponential dependence on the unknown parameter’s norm, which was a fundamental limitation in most previous works. Moreover, we establish a near-matching lower bound of $\Omega\Big(\frac{d}{K\sqrt{T}}\Big)$, where $K$ is the maximum subset size. To the best of our knowledge, this is the first theoretical result in PbRL with ranking feedback that explicitly shows improved sample efficiency as a function of the subset size.
@@ 1 Introduction
The framework of Preference-based Reinforcement Learning (PbRL) [busa2014preference, wirth2016model, wirth2017survey, saha2023dueling] was introduced to address the difficulty of designing effective reward functions, which often demands substantial and complex engineering effort [wirth2013preference, wirth2017survey]. PbRL has been successfully applied in diverse domains, including robot training, stock prediction, recommender systems, and clinical trials [jain2013learning, sadigh2017active, christiano2017deep, kupcsik2017learning, novoseller2020dueling]. Notably, PbRL also serves as a foundational framework for Reinforcement Learning from Human Feedback (RLHF) when feedback is provided in the form of preferences rather than explicit scalar rewards. This preference-based approach has proven highly effective in aligning Large Language Models (LLMs) with human values and preferences [christiano2017deep, ouyang2022training, rafailov2023direct].
@@
Given its practical success, the field has also seen significant theoretical advances [chen2022human, mehta2023sample, saha2023dueling, zhu2023principled, xiong2023iterative, zhan2023provable, zhan2023provableoffline, wu2023making, sekhari2023contextual, munos2023nash, cen2024value, chowdhury2024provably, dong2024rlhf, das2024active, mukherjee2024optimal, swamy2024minimaximalist, scheid2024optimal, thekumparampil2024comparing, xie2024exploratory, chen2025avoiding, kveton2025active]. However, despite this progress, most existing models remain limited to handling only pairwise comparison feedback. A few works [zhu2023principled, mukherjee2024optimal, thekumparampil2024comparing] explore the more general setting of multiple comparisons, offering a strict extension beyond the pairwise case. zhu2023principled study the offline setting, where a dataset of questions (or contexts) along with corresponding ranking feedback over $K$ answers (or actions), labeled by human annotators, is available. mukherjee2024optimal investigate the online learning-to-rank problem [radlinski2008learning], where a dataset of questions with $K$ candidate answers is provided, but no feedback is initially available. thekumparampil2024comparing consider a context-free setting (i.e., a singleton context), and the goal is to learn the ranking of $N\geq K$ answers based on ranking feedback obtained from subsets of size $K$. However, all of their theoretical performance guarantees fail to show that using multiple comparisons provides any advantage over the pairwise setting (see Table 1). This is counterintuitive, as ranking feedback is inherently more informative than pairwise feedback. Specifically, since a ranking over $K$ actions provides $\binom{K}{2}$ pairwise comparisons, it should, in principle, enable faster learning and lead to stronger performance guarantees.
@@
Table 1: Comparisons of settings and theoretical guarantees in related works on PbRL with ranking feedback. Here, $T$ denotes the number of rounds (or the number of data points in the offline setting), $K$ is the (maximum) size of the offered action set (i.e., assortment), and $d$ is the feature dimension. $\rho$ represents the unknown context distribution. Here, $\tilde{\mathcal{O}}$ hides logarithmic factors and polynomial dependencies on $B$. “Pred. Error” refers to the prediction error.
@@
zhu2023principled Offline Accessible $\mathcal{X}$ Given Suboptimality $\tilde{\mathcal{O}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}e^{B}K^{2}}\sqrt{\frac{d}{T}}\right)$
mukherjee2024optimal Online Accessible $\mathcal{X}$ Given Pred. Error $\tilde{\mathcal{O}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}e^{B}K^{3}}\frac{d}{\sqrt{T}}\right)$
thekumparampil2024comparing Online No context Select $K$ Pred. Error $\tilde{\mathcal{O}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}e^{B}K^{3}}\frac{d}{\sqrt{T}}\right)$
@@
In this paper, we assume that the ranking feedback follows the Plackett-Luce (PL) model [plackett1975analysis, luce1959individual], where, in each round, the learner receives ranking feedback over a subset of up to $K$ actions (with $K\leq N$) selected from a universe of $N$ actions. This problem setup is closely related to that of thekumparampil2024comparing; however, unlike their work, which focuses solely on a context-free setting (or equivalently, a fixed singleton context), we study a more general setting where contexts are diverse and drawn from an unknown distribution.
@@
Furthermore, our suboptimality gap eliminates the exponential dependence on the parameter norm bound, $\mathcal{O}(e^{B})$, in the leading term. This improvement stems from analytically dividing the total rounds into warm-up and non–warm-up phases (see the proof sketch in Section 5.1). This represents a significant improvement over most prior works, where performance guarantees depend on $\mathcal{O}(e^{B})$ [saha2021optimal, saha2023dueling, zhu2023principled, xiong2023iterative, zhan2023provable, das2024active, xie2024exploratory, thekumparampil2024comparing, kveton2025active]. Very recently, a few works [chen2025avoiding, di2024nearly] have successfully avoided the $\mathcal{O}(e^{B})$ dependency by relying on auxiliary techniques or additional information—such as specialized sampling schemes [chen2025avoiding] or prior knowledge of $\kappa$ [di2024nearly]—which, however, are often impractical. Moreover, their methods are limited to pairwise comparison settings. In contrast, our approach eliminates the $\mathcal{O}(e^{B})$ dependency without using any auxiliary techniques and considers more general ranking feedback beyond pairwise comparisons. Our main contributions are summarized as follows:
@@
  • Improvement on $\mathcal{O}(e^{B})$ dependency: Our result eliminates the exponential dependence on the parameter norm bound, $\mathcal{O}(e^{B})$, in the leading term—without relying on any auxiliary techniques—through a fully refined analysis. This shows that the $\mathcal{O}(e^{B})$ dependence widely observed in PbRL (or RLHF) and dueling bandit analyses is not inherent but rather an artifact of loose analysis. Moreover, our key technique for removing the $\mathcal{O}(e^{B})$ dependence—by dividing the rounds into warm-up and non–warm-up phases only in the analysis—can be seamlessly incorporated into existing PbRL or dueling bandit analyses, including regret-minimization frameworks without altering the underlying algorithms (see Appendix G.1). To the best of our knowledge, this is the first PbRL work with ranking feedback beyond pairwise comparisons that avoids the $\mathcal{O}(e^{B})$ dependence.
  • Efficiency of rank-breaking (RB): We present both naive PL loss–based and rank-breaking (RB) loss–based learning approaches, and show that the RB formulation achieves superior computational efficiency and empirical performance compared to the PL counterpart.
  • Lower bound: We establish a near-matching lower bound of $\Omega\!\left(\frac{d}{K\sqrt{T}}\right)$ under the PL model with ranking feedback, matching our upper bound up to a $K$ factor. This shows that leveraging richer ranking information (larger $K$) provably improves sample efficiency.
@@ 2 Related Works
Fueled by the remarkable success of LLMs [christiano2017deep, ouyang2022training, rafailov2023direct], the theoretical study of PbRL has rapidly emerged as a central focus within the research community. Early work in this area traces back to the dueling bandits literature [yue2012k, zoghi2015copeland, saha2018battle, bengs2021preference].
@@
Dueling bandits. The dueling bandit framework, introduced by yue2012k, departs from the classical multi-armed bandit setting by requiring the learner to select two arms and observe only their pairwise preference. For general preferences, a single best arm that is globally dominant may not exist. To address this, various alternative winners have been proposed, including the Condorcet winner [zoghi2014relative, komiyama2015regret], Copeland winner [zoghi2015copeland, wu2016double, komiyama2016copeland], Borda winner [jamieson2015sparse, falahatgar2017maxing, heckel2018approximate, saha2021adversarial, wu2023borda], and von Neumann winner [ramamohan2016dueling, dudik2015contextual, balsubramani2016instance], each with its own corresponding performance metric.
@@
To address scalability and contextual information, saha2021optimal proposed a structured contextual dueling bandit setting in which preferences are modeled using a Bradley–Terry–Luce (BTL) model [bradley1952rank] based on the unknown intrinsic rewards of each arm. In a similar setting, bengs2022stochastic studied a contextual linear stochastic transitivity model, and di2023variance proposed a layered algorithm that achieves variance-aware regret bounds. However, most prior dueling bandit works suffer from an $\mathcal{O}(e^{B})$ dependence. Only a few recent studies [di2024nearly, chen2025avoiding] have succeeded in removing this $\mathcal{O}(e^{B})$ term, either by introducing additional complex subroutines [chen2025avoiding] or by relying on prior knowledge of $\kappa$ [di2024nearly].
@@
Preference-based reinforcement learning (PbRL). Building upon this line of work, subsequent research has extended the dueling bandit framework to the RL, considering both online [xu2020preference, novoseller2020dueling, chen2022human, saha2023dueling, wu2023making] and offline settings [zhu2023principled, zhan2023provable, liu2024provably]. More recently, under the active learning framework—where the full set of contexts $\mathcal{X}$ is accessible—many studies aim to improve sample efficiency by selecting prompts either based on the differences in estimated rewards for their responses [muldrew2024active] or through D-optimal design methods [mehta2023sample, scheid2024optimal, das2024active, mukherjee2024optimal, thekumparampil2024comparing, kveton2025active]. However, most of these works focus exclusively on pairwise preference feedback and cannot be extended to more general ranking feedback cases. mukherjee2024optimal study the online learning-to-rank problem when prompts are given along with $K$ candidate answers, while thekumparampil2024comparing investigate learning to rank $N\geq K$ answers from partial rankings over $K$ answers, but under a context-free setting. In this paper, we consider a stochastic contextual setting (more general than thekumparampil2024comparing), where contexts are sampled from an unknown but fixed distribution, and aim to minimize the suboptimality gap using ranking feedback of up to length $K$. For further related work, see Appendix A.
@@ 3 Problem Setting and Preliminaries
Let ${\mathcal{S}}:=\{S\subseteq\mathcal{A}\mid 2\leq|S|\leq K\}$ be the collection of all action subsets whose sizes range from $2$ to $K$. For any $S\in{\mathcal{S}}$, let $\sigma$ denote the labeler’s ranking feedback—that is, a permutation of the elements in $S$. We write $\sigma_{j}$ for the $j$-th most preferred action under $\sigma$. We model the distribution of such rankings using the Plackett-Luce (PL) model [plackett1975analysis, luce1959individual], defined as:
@@
Here, $r_{\bm{\theta}^{\star}}$ represents a reward model parameterized by the unknown parameter $\bm{\theta}^{\star}$.
When K=2K\!=\!2, this reduces to the pairwise comparison framework considered in the Bradley-Terry-Luce (BTL) model [bradley1952rank]. The probability that $a$ is preferred to $a^{\prime}$ given $x$ can be expressed as:
@@
Let $\phi:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ be a known feature map satisfying $\max_{x,a}\|\phi(x,a)\|_{2}\leq 1$, and let $\bm{\theta}^{\star}\in\mathbb{R}^{d}$ denote the true but unknown parameter. The reward is assumed to follow a linear structure given by $r_{\bm{\theta}^{\star}}(x,a)=\phi(x,a)^{\top}\bm{\theta}^{\star}$. We further assume realizability, i.e., $\bm{\theta}^{\star}\in\Theta:=\{\bm{\theta}\in\mathbb{R}^{d}\mid\|\bm{\theta}\|_{2}\leq B\}$. Without loss of generality, we assume $B\geq 1$.
At each round $t\in[T]$, a context $x_{t}\in\mathcal{X}$ is drawn from a fixed but unknown distribution $\rho$. Given the context $x_{t}$, the learning agent selects a subset of actions $S_{t}\in{\mathcal{S}}$—referred to as an assortment throughout the paper—and receives a ranking over $S_{t}$ as feedback, generated according to the PL model. Let $\pi^{\star}(x)=\mathop{\mathrm{argmax}}_{a}r_{\bm{\theta}^{\star}}(x,a)$ be the optimal policy under the true reward $r_{\bm{\theta}^{\star}}$. After $T$ rounds of interaction with the labeler, the goal is to output a policy $\widehat{\pi}_{T}:\mathcal{X}\rightarrow\mathcal{A}$ that minimizes the suboptimality gap, defined as:
$\textup{{SubOpt}}(T):=\mathbb{E}_{x\sim\rho}\left[r_{\bm{\theta}^{\star}}\left(x,\pi^{\star}(x)\right)-r_{\bm{\theta}^{\star}}\left(x,\widehat{\pi}_{T}(x)\right)\right].$
3.1 Loss Functions and Rank-Breaking